OpenStudy (anonymous):

2*n+2*n-1 divided by 2n+1-2*n= 3/2

mathslover (mathslover):

is N and n same?

OpenStudy (anonymous):

n

mathslover (mathslover):

ok .. \[\large{\frac{2n+2n-1}{2n-1-2n}=\frac{3}{2}}\]

mathslover (mathslover):

\[\large{4n-1=\frac{6n-3-6n}{2}}\] \[\large{8n-2=6n-3-6n}\]

OpenStudy (anonymous):

yes

mathslover (mathslover):

\[\large{8n-2=-3}\] \[\large{8n-2+3=0}\]

mathslover (mathslover):

can you solve further ? ?

mathslover (mathslover):

@shivraj .. what r u getting for the value for n ?

OpenStudy (anonymous):

no i dont hav any idea

mathslover (mathslover):

ok .. : -2 + 3 =?

OpenStudy (anonymous):

1

mathslover (mathslover):

right .. now 8n -1 = 0 : can u do it now?

OpenStudy (anonymous):

yes

mathslover (mathslover):

so what r u getting?

OpenStudy (anonymous):

3/2

mathslover (mathslover):

8n - 1 = 0 8n - 1 + 1 = 0 +1 8n = 1 n = 1/8 ..

mathslover (mathslover):

check it again @shivraj

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is 1/8 the answer @mathslover

OpenStudy (anonymous):

See, here: \[\large 8n - 2 + 3 = 0 \implies 8n = -1\] \[\large \color{green}{n = \frac{-1}{8}}\]

OpenStudy (anonymous):

@shivraj are you getting??

mathslover (mathslover):

oh yes it is -1/8 ..

OpenStudy (anonymous):

thnks